Grade 12th passPhysical Chemistry

Fe(oh)2 is diacidic base has kb1=(10)-4 and kb2=2.5*(10)-6. What Is the concentration of fe(oh)2 in 0.1M fe(no3)2 soln.

Jit
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago

kb1=(10)- 4 and kb2=2.5*(10)- 6 
Convergence of ferrous nitrate = 0.1 M 
At that point convergence of ferrous hydroxide = ? 
Kb = ([HB+] [OH¯])/[B] 
10 - 4 = 0.1 x 2.5* 10 ^ - 6 concn 
10-4=0.25 * 10^6 conc 
Concn = 0.25 * 10^6/10-4 
Concn = 0.0005mole/l